2/x+2 + 3/x-2 = 5/X^2-4
Answer by jorel1380(3719)

2/(x+2)+3/(x-2)=5/(x2-4)
2/(x+2)+3/(x-2)=5/(x+2)(x-2)
2(x-2)+3(x+2)=5
2x-4+3x+6=5
5x+2=5
5x=3
x=3/5..